TORSIONAL ANALYSIS OF 2-BUTYNOL

\underline{RANGANATHAN SUBRAMANIAN}, Department of Chemistry, Wesleyan University, Middletown, CT 06459; ROBERT K. BOHN, Department of Chemistry, University of Connecticut, Storrs, CT 06269; and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

Microwave spectroscopic studies of substituted propynes, CH_{3}-C≡C-X, have increased our understanding of the nature and origin of the potential barriers for internal rotation. Specifically, the potential barriers have been measured for the groups, -CD_{3}^{a}, -CH_{2}Cl^{b}, -SiH_{3}^{c}, and -COF^{d}, bonded to the propynyl fragment, with reported 3-fold barriers of 5.62(16), 10.05(9), 3.77(70) and 2.20(12) cm^{−1} respectively. We report here a study of the rotational spectrum of 2-butynol, CH_{3}-C≡C-CH_{2}OH. 2-butynol has two low energy rotational conformers with the OH group either anti or gauche with respect to the propynyl moiety. The study of 2-butynol was performed with a pulsed-jet, Fourier Transform Microwave Spectrometer. The rotational spectrum of one species including tunneling splittings has been assigned and is consistent with the gauche conformer. A preliminary value of the 3-fold barrier is 12.5(4) cm^{−1}. Unassigned transitions in the spectrum may be due to the other conformer.

aJ. Nakagawa, M. Hayashi, Y. Endo, S. Saito, and E. Hirota J. Chem. Phys. 80, 5922 (1966)

bV.M. Stolwijk and B.P. van Eijck J. Mol. Spectroc. 124, 92 (1987)

cJ. Nakagawa, K. Yamada, M. Bester, and G. Winnewisser J. Mol. Spectrosc. 110, 74 (1985)

dK. D. Hensel and M. C. L. Gerry J. Chem. Soc. Faraday Trans. 90, 3023 (1992)